The distribution of weight measurements of adult women in Neverland is approximately Normal with mean 166 pounds and standard deviation 20 pounds.

  1. What percentage of the population is between 1SD and 2SD below the mean? (2p)
Answer #1

Solution :

166 - 20 = 146 and 166 - 40 = 126

P(126 < x < 146) = P[(126 - 166)/ 20) < (x - ) /  < (146 - 166) / 20) ]

= P(-2 < z < -1)

= P(z < -1) - P(z < -2)

= 0.1587 - 0.0228

= 0.1359

Answer = 13.59%